OpenStudy (anonymous):

how do you solve 8/6=x/15

OpenStudy (anonymous):

you must get the X on one side of the equation, or for your example 8/6=x/15 you could cross multiply. so 8*15 = 6*X and then get x on one side of the equation there so 8*15 over 6 = X

OpenStudy (anonymous):

8/6=x/15<=>8*15=6*x<=>x=8*15/6<=>x=120/6<=>x=20